1. A tank contains 200 liters of fluid in which 20 grams of salt is dissolved. Brine containing 1 gram of salt per liter is then pumped into the tank at a rate of 5 L/min; the well-mixed solution is pumped out at the same rate. Find the number A(t) of grams of salt in the tank at time t.

A(t) = ____ g

2. A tank contains 420 liters of fluid in which 30 grams of salt is dissolved. Pure water is then pumped into the tank at a rate of 6 L/min; the well-mixed solution is pumped out at the same rate. Find the number A(t) of grams of salt in the tank at time t.

A(t) = ___ g
1
Expert's answer
2020-06-22T17:26:44-0400

(1)


initial amount of salt = 20 gram


rate of pumping in = 5L/min


Let A be amount of salt at any time t,

then


"\frac{dA}{dt} = R_{in} - R_{out}"


R is rate of flow.


"R_{in} = \frac{1g}{1L} \frac{5L}{1 min} = 5 g/min"


"R_{out} = \frac{A}{200}*{5} = \frac{A}{40} g/min"


Hence equation of state will be,


"\frac{dA}{dt} = 5-\frac{A}{40}"


solving equation,


"\frac{dA}{dt} + \frac{1}{40}A =5"



"e^{\frac{1}{40}t}A = \int 5e^{\frac{1}{40}t}dt"


"e^{\frac{1}{40}t}A = 200e^{\frac{1}{40}t} + C"


"A = 200 + Ce^{-\frac{1}{40}t}"


applying condition, A=20 at t=0


"20 = 200 + C \implies C = -180"


Hence desired equation of the state is given by,


"A = 200 - 180e^{-\frac{1}{40}t}"




(2) For this part, since we are pumping pure water so "R_{in} = 0"


similarly, making the equation of the state,


"\frac{dA}{dt} = R_{in} - R_{out}"


"R_{out} = \frac{A}{420}*6 = \frac{A}{70} g/min"


Then


"\frac{dA}{dt} = -\frac{A}{70}"


solving equation,


"\int \frac{dA} {A} = - \frac{1}{70} \int dt"


"ln A = -\frac{1}{70}t + C"


Applying condition that at t=0, A = 30


"ln(30) = C"


then


"lnA = -\frac{1}{70}t +ln(30)"


"A = 30e^{-\frac{1}{70}t}"